# Neural Posterior Estimation for Cataloging Astronomical Images from the Legacy Survey of Space and Time

### v1

@@ Abstract
show]cavestruz@umich.edu
show]regier@umich.edu
@@ 1 Introduction
The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) will begin full-scale operation in 2026. During its 10-year lifespan, LSST will collect 500 petabytes of images of the night sky in the southern hemisphere. The survey aims to address questions about the structure and evolution of the Universe pertaining to the nature of dark energy and dark matter, the contents of the solar system and the transient optical sky, and the structure of the Milky Way (Ž. Ivezić et al., 2019).
@@
Inferring the correct catalog for imaged light sources that visually overlap, known as blended objects or blends, is even more challenging; it requires separating each light source from both background noise and other light sources. As advances in telescope technology improve light sensitivity, resulting in a higher density of light sources in the images captured, blending has become increasingly common (P. Melchior et al., 2021). LSST will observe a much higher density of sources than previous wide-field surveys. Projections indicate that 62% of the LSST-imaged galaxies will be blended to some extent (J. Sanchez et al., 2021), which is in line with figures for the Hyper Suprime-Cam survey, in which 58% of objects are blended (J. Bosch et al., 2017).
@@
Traditional algorithmic approaches to cataloging, such as the approaches of SExtractor (E. Bertin & S. Arnouts, 1996) and the planned LSST image processing pipeline (J. Bosch et al., 2018), do not address the ambiguity posed by the increased volume and enhanced resolution of LSST images in a statistically coherent way. For instance, in the LSST pipeline, detections are made independently for each image band using a peak-finding routine. Following this, detections across bands are merged according to a matching criterion. Blended objects are often subject to over-merging during this stage, so candidate blends are subsequently processed by a deblending algorithm. The primary deblending algorithm employed by LSST, known as Scarlet, operates under the assumption that light sources are compact and their spectra are spatially uniform (P. Melchior et al., 2018). Despite the inherent ambiguity in detecting and characterizing light sources, traditional algorithms ultimately produce a single realization of the catalog.
@@
Probabilistic cataloging is an alternative to traditional algorithmic approaches (B. J. Brewer et al., 2013; S. K. Portillo et al., 2017; R. M. Feder et al., 2020; J. J. Buchanan et al., 2023). In probabilistic cataloging, a distribution of latent catalogs is inferred conditional on the images, which are treated as observed random variables in a generative model. Deblending is a consequence of performing posterior inference, eliminating the need for special process logic to address potential blends. In addition, probabilistic cataloging effectively handles multiband imaging, in which each image captures photons in different frequency ranges, by modeling each image as a realization of the same latent catalog (R. M. Feder et al., 2020).
@@
Unfortunately, existing probabilistic cataloging methods cannot be applied to the complete LSST dataset. Probabilistic cataloging methods based on Markov chain Monte Carlo (MCMC), including B. J. Brewer et al. (2013); S. K. Portillo et al. (2017); R. M. Feder et al. (2020) and J. J. Buchanan et al. (2023), are not computationally efficient enough to apply to the entire collection of LSST images; these methods are designed to process small cutouts of much larger images. Additionally, B. J. Brewer et al. (2013); S. K. Portillo et al. (2017); R. M. Feder et al. (2020) and R. Liu et al. (2023) are limited to detecting only stars, rather than both stars and galaxies. J. Regier et al. (2019) refines the estimates of preexisting catalogs, but is unable to identify new stars and galaxies or account for detection ambiguity.
@@
In this work, we explore an approach to Bayesian inference known as neural posterior estimation (NPE). NPE is a type of likelihood-free amortized variational inference that fits a variational distribution to minimize an expectation of the Kullback-Leibler (KL) divergence from the posterior to the variational distribution (K. Zhang et al., 2021). Algorithmically, NPE amounts to simulating data (catalogs and images) from the generative model and fitting a neural network (the inference network) to predict the corresponding catalog. Recently, NPE has been shown to be globally convergent, with plausible assumptions, despite the non-convexity of neural networks with respect to their parameters (D. McNamara et al., 2024). Most importantly, for coadded images, NPE implicitly marginalizes nuisance latent variables (L. Ambrogioni et al., 2019). Thus, despite performing inference with a properly specified Bayesian model that incorporates an intractable quantity of nuisance latent variables, NPE remains computationally efficient.
@@
In our collaboration’s previous work with NPE for probabilistic cataloging (R. Liu et al., 2023; D. Hansen et al., 2022; A. Patel et al., 2025; J. Regier, 2025), we performed inference exclusively with single-exposure images, and in a workshop paper, we put forward the idea of using NPE for coadd images (M. Wang et al., 2022). In this work, we apply an NPE method (Section 3) to images that are coadded and multiband. To the best of our knowledge, this is the first archival publication to report applying Bayesian inference to coadded astronomical images with a model that accounts for the dependence among pixels due to the coaddition process.
@@ 2 The LSST DESC DC2 Simulated Sky Survey
In preparation for the arrival of LSST images, the LSST Dark Energy Science Collaboration created the DC2 dataset as a testbed for image-processing software currently in development (D. Korytov et al., 2019; B. Abolfathi et al., 2021a, b). The dataset has three components: a large synthetic galaxy catalog designed for dark energy science (Section 2.1); a collection of simulated unprocessed images, resembling the images read off a CCD (Section 2.2); and a collection of calibrated, coadded, and otherwise processed images (Section 2.3). The DC2 simulator effectively serves as the generative model for the proposed method (Section 2.4) as well as providing our validation data.
@@ 2.1 Catalog simulation
CosmoDC2, introduced by D. Korytov et al. (2019), is a sophisticated synthetic galaxy catalog developed to aid dark energy science analyses, specifically aimed at constraining cosmological parameters. The CosmoDC2 simulation incorporates three inputs:
@@
    The Outer Rim simulation (K. Heitmann et al., 2019), a trillion-particle n-body simulation. Outer Rim is one of the largest simulations at its mass resolution and was run using the Hardware/Hybrid Accelerated Cosmology Code (S. Habib et al., 2016).
@@
    The Universe Machine catalog (P. Behroozi et al., 2019), which contains the output of an n-body simulation that models the connection between galaxies and dark matter halos. In the simulation, galaxy positions, stellar masses, and star formation rates are empirically calibrated with observational data.
@@
    The Galacticus library is based on a semi-analytic model of galaxy formation (A. J. Benson, 2012) applied to the AlphaQ simulation. The semi-analytic model maps galaxy evolution to the dark matter halo mergers of the simulation.
@@ 2.2 Simulation of raw images
The simulated images include six optical bands, each composed of $4000\times 4000$ pixels. The images are simulated using GalSim (B. Rowe et al., 2015) in conjunction with LSST-specific GalSim extensions provided by the imSim package. The imSim package, in turn, is based on the Rubin Observatory image simulation package, which provides a simulator of the LSST camera (LSSTCam) that models its 189 science charge-coupled devices (CCDs), each with 16 megapixels. The imSim package models the characteristics of the LSST filter, the brightness of the sky (background), and image artifacts, such as cosmic rays, pixel saturation, and bleed trails.
@@ 3 Methodology: Bayesian inference
The method development in this section follows that of J. Regier (2025), which introduces a spatially autoregressive variational distribution for the detection of astronomical objects, and establishes its theoretical and empirical properties. Here, we generalize this variational distribution to apply to multiband coadded images.
@@ 3.2 Our variational family
J. Regier (2025) establishes conditions under which a variational distribution of this form mirrors the conditional independence assertions of the target distribution $p(z\mid x)$, in the sense of being a “minimally sufficient I-map.” This is a desirable property for a variational distribution because it implies that the variational distribution has the flexibility to model dependencies in the target distribution, but not enough flexibility to model dependencies that are known a priori to be absent in the target distribution.
@@ 3.3 Neural Posterior Estimation
During training, we additionally regularize by performing on-the-fly data augmentation. We follow standard practice in neural network performance optimization, which involves using validation loss to search for good network architectures and tuning parameters (A. Karpathy, 2019). Training is computationally intensive, requiring 22 hours on a single GPU. Inference is extremely fast, with data loading (moving images from disk to CPU memory) acting as the bottleneck rather than the neural network forward passes required to process that data. Appendix B provides further details of the proposed algorithm, including illustrations of the neural network architecture, projected runtimes for our methods on the complete LSST coadd image set, and discussion about storing probabilistic LSST catalogs.
@@ 4.1 Our primary benchmark: the LSST coadd catalog
The LSST pipeline constructs catalogs of objects in coadded images through a sequential multi-step process (J. Bosch et al., 2017).
@@ 4.2 Matching detections and true objects
To evaluate the accuracy of a catalog, we first match cataloged detections and true objects. The true objects available for matching are those exceeding 50 Jy in flux. For each of these objects, we identify the corresponding detection by selecting the nearest detected object within a one-pixel radius. A one-pixel radius is within the range of matching thresholds used in other publications: J. Sánchez et al. (2020) reports using a three-pixel radius for DC1 data and S. K. Portillo et al. (2017) reports using a half-pixel radius for SDSS data. Appendix C considers additional matching thresholds, namely three and five pixels; in short, the general trends in relative performance are not changed by the precise threshold used. Among detection-object pairs within the specified radius, we then find the optimal (distance-minimizing) bipartite matching using the Hungarian algorithm. The DC2 catalog provides a matching based on a different matching algorithm, but we exclusively use our matching procedure to ensure that the matching is done in the same way for both BLISS and the LSST pipeline.
@@ 4.6 Units of flux
The flux (or brightness) of an object directly influences the accuracy with which a method can detect that object. We quantify the brightness of an object in units of AB magnitude (Ž. Ivezić & the LSST Project Science Team, 2018). In this system, magnitude is defined as $2.5\log_{10}(F_{b}/F_{AB})$, where $F_{b}$ is the flux of an object in Jy (Jansky) and $F_{AB}$ is the zero point, which is set to 3631 Jy. Note that smaller magnitudes correspond to brighter sources.
@@ 4.8 Units of blending
We report some metrics stratified by blendedness, typically in the context of a minimum flux threshold. We adopt the high-level approach to quantifying blendedness of J. Bosch et al. (2017), in which blendedness is $1-\frac{I_{\textrm{child}}}{I_{\textrm{parent}}}$, where $I_{\textrm{child}}$ is the flux of the object in question (true or predicted, depending on context) and $I_{\textrm{parent}}$ is the overall flux in the vicinity of this object. The closer this value is to 1, the more blended the object is.
We define $I_{\textrm{child}}$ and $I_{\textrm{parent}}$ slightly differently than J. Bosch et al. (2017) because we alone have access to ground-truth catalogs by virtue of working with synthetic data. We take $I_{\textrm{child}}$ to be the true flux of the object in question (that is, without shot noise) and we take $I_{\textrm{parent}}$ to be the Gaussian-weighted sum of the true fluxes of neighboring objects. The covariance is isotropic with a standard deviation of 6 pixels (1.2 arc seconds), and we truncate beyond 30 pixels (6 arc seconds).
@@ 5 Experimental Results
In this section, we refer to our implementation of the proposed NPE method as the Bayesian Light Source Separator (BLISS). BLISS is a long-running software project (R. Liu et al., 2023; D. Hansen et al., 2022; A. Patel et al., 2025; J. Regier, 2025), which we have extended in this work to apply to the DC2 dataset. Our experiments use the benchmark methodology from Section 4 to assess the performance of BLISS in object detection, flux estimation, star/galaxy classification, and galaxy shape estimation.
@@ 5.1 Detection
To generalize the concepts of precision and recall to spatial data, in keeping with S. K. Portillo et al. (2017) and others, we define detection precision (purity) as the proportion of detections that are successfully matched with true objects and detection recall (completeness) as the proportion of objects that are successfully matched with detections. Figure 3 presents detection precision and recall as a function of magnitude for both BLISS and the LSST pipeline. The histogram in the top-right panel of Figure 3 shows the total number of true objects in each r-band magnitude bin. Compared to LSST, BLISS successfully identifies a larger fraction of sources in each of the magnitude bins. Within the r-band magnitude range of 23.9 to 25.6, the recall of LSST decreases from 93% to 58%, and the recall of BLISS exhibits a similar drop from 97% to 64%. The difference in recall between LSST and BLISS is largest in the second-to-last magnitude bin, where the difference in recall exceeds 10%. Even for the faintest sources (magnitude >> 25.6), which are the most difficult to detect, BLISS outperforms the LSST pipeline in terms of recall by 8%.
@@ 5.3 Star/galaxy classification
Classifying stars and galaxies in photometric surveys is essential for various scientific applications, including studies of galaxy formation and cosmology (I. Sevilla-Noarbe et al., 2018). In the field of weak lensing, identifying a pure sample of stars is crucial for fitting a model of the point spread function for use in inferring cosmic shear. Mistaking galaxies for stars could result in spurious correlations.
Star/galaxy separation is challenging because many light sources are faint and blended. Further, once convolved with the PSF, some galaxies appear scarcely larger than stars. The LSST pipeline does not use machine learning to separate stars and galaxies. Instead, it compares the fit of an extended “composite” model to that of a point source (J. Bosch et al., 2018)

### v2

@@ Abstract
[…]
@@ 1 Introduction
The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) will begin full-scale operation in 2026. During its 10-year lifespan, LSST will collect 500 petabytes of images of the night sky in the southern hemisphere. The survey aims to address questions about the structure and evolution of the Universe pertaining to the nature of dark energy and dark matter, the contents of the solar system and the transient optical sky, and the structure of the Milky Way (Ivezić et al., 2019).
@@
Inferring the correct catalog for imaged light sources that visually overlap, known as blended objects or blends, is even more challenging; it requires separating each light source from both background noise and other light sources. As advances in telescope technology improve light sensitivity, resulting in a higher density of light sources in the images captured, blending has become increasingly common (Melchior et al., 2021). LSST will observe a much higher density of sources than previous wide-field surveys. Projections indicate that 62% of the LSST-imaged galaxies will be blended to some extent (Sanchez et al., 2021), which is in line with figures for the Hyper Suprime-Cam survey, in which 58% of objects are blended (Bosch et al., 2017).
@@
Traditional algorithmic approaches to cataloging, such as the approaches of SExtractor (Bertin & Arnouts, 1996) and the planned LSST image processing pipeline (Bosch et al., 2018), do not address the ambiguity posed by the increased volume and enhanced resolution of LSST images in a statistically coherent way. For instance, in the LSST pipeline, detections are made independently for each image band using a peak-finding routine. Following this, detections across bands are merged according to a matching criterion. Blended objects are often subject to over-merging during this stage, so candidate blends are subsequently processed by a deblending algorithm. The primary deblending algorithm employed by LSST, known as Scarlet, operates under the assumption that light sources are compact and their spectra are spatially uniform (Melchior et al., 2018). Despite the inherent ambiguity in detecting and characterizing light sources, traditional algorithms ultimately produce a single realization of the catalog.
@@
Probabilistic cataloging is an alternative to traditional algorithmic approaches (Brewer et al., 2013; Portillo et al., 2017; Feder et al., 2020; Buchanan et al., 2023). In probabilistic cataloging, a distribution of latent catalogs is inferred conditional on the images, which are treated as observed random variables in a generative model. Deblending is a consequence of performing posterior inference, eliminating the need for special process logic to address potential blends. In addition, probabilistic cataloging effectively handles multiband imaging, in which each image captures photons in different frequency ranges, by modeling each image as a realization of the same latent catalog (Feder et al., 2020).
@@
Unfortunately, existing probabilistic cataloging methods cannot be applied to the complete LSST dataset. Probabilistic cataloging methods based on Markov chain Monte Carlo (MCMC), including Brewer et al. (2013); Portillo et al. (2017); Feder et al. (2020) and Buchanan et al. (2023), are not computationally efficient enough to apply to the entire collection of LSST images; these methods are designed to process small cutouts of much larger images. Additionally, Brewer et al. (2013); Portillo et al. (2017); Feder et al. (2020) and Liu et al. (2023) are limited to detecting only stars, rather than both stars and galaxies. Regier et al. (2019) refines the estimates of preexisting catalogs, but is unable to identify new stars and galaxies or account for detection ambiguity.
@@
In this work, we explore an approach to Bayesian inference known as neural posterior estimation (NPE). NPE is a type of likelihood-free amortized variational inference that fits a variational distribution to minimize an expectation of the Kullback-Leibler (KL) divergence from the posterior to the variational distribution (Zhang et al., 2021). Algorithmically, NPE amounts to simulating data (catalogs and images) from the generative model and fitting a neural network (the inference network) to predict the corresponding catalog. Recently, NPE has been shown to be globally convergent, with plausible assumptions, despite the non-convexity of neural networks with respect to their parameters (McNamara et al., 2024). Most importantly, for coadded images, NPE implicitly marginalizes nuisance latent variables (Ambrogioni et al., 2019). Thus, despite performing inference with a properly specified Bayesian model that incorporates an intractable quantity of nuisance latent variables, NPE remains computationally efficient.
@@
In our collaboration’s previous work with NPE for probabilistic cataloging (Liu et al., 2023; Hansen et al., 2022; Patel et al., 2025; Regier, 2025), we performed inference exclusively with single-exposure images, and in a workshop paper, we put forward the idea of using NPE for coadd images (Wang et al., 2022). In this work, we apply an NPE method (Section 3) to images that are coadded and multiband. To the best of our knowledge, this is the first archival publication to report applying Bayesian inference to coadded astronomical images with a model that accounts for the dependence among pixels due to the coaddition process.
@@ 2 The LSST DESC DC2 Simulated Sky Survey
In preparation for the arrival of LSST images, the LSST Dark Energy Science Collaboration created the DC2 dataset as a testbed for image-processing software currently in development (Korytov et al., 2019; Abolfathi et al., 2021a, b). The dataset has three components: a large synthetic galaxy catalog designed for dark energy science (Section 2.1); a collection of simulated unprocessed images, resembling the images read off a CCD (Section 2.2); and a collection of calibrated, coadded, and otherwise processed images (Section 2.3). The DC2 simulator effectively serves as the generative model for the proposed method (Section 2.4) as well as providing our validation data.
@@ 2.1 Catalog simulation
CosmoDC2, introduced by Korytov et al. (2019), is a sophisticated synthetic galaxy catalog developed to aid dark energy science analyses, specifically aimed at constraining cosmological parameters. The CosmoDC2 simulation incorporates three inputs:
@@
    The Outer Rim simulation (Heitmann et al., 2019), a trillion-particle n-body simulation. Outer Rim is one of the largest simulations at its mass resolution and was run using the Hardware/Hybrid Accelerated Cosmology Code (Habib et al., 2016).
@@
    The Universe Machine catalog (Behroozi et al., 2019), which contains the output of an n-body simulation that models the connection between galaxies and dark matter halos. In the simulation, galaxy positions, stellar masses, and star formation rates are empirically calibrated with observational data.
@@
    The Galacticus library is based on a semi-analytic model of galaxy formation (Benson, 2012) applied to the AlphaQ simulation. The semi-analytic model maps galaxy evolution to the dark matter halo mergers of the simulation.
@@ 2.2 Simulation of raw images
The simulated images include six optical bands, each composed of $4000\times 4000$ pixels. The images are simulated using GalSim (Rowe et al., 2015) in conjunction with LSST-specific GalSim extensions provided by the imSim package. The imSim package, in turn, is based on the Rubin Observatory image simulation package, which provides a simulator of the LSST camera (LSSTCam) that models its 189 science charge-coupled devices (CCDs), each with 16 megapixels. The imSim package models the characteristics of the LSST filter, the brightness of the sky (background), and image artifacts, such as cosmic rays, pixel saturation, and bleed trails.
@@ 3 Methodology: Bayesian inference
The method development in this section follows that of Regier (2025), which introduces a spatially autoregressive variational distribution for the detection of astronomical objects, and establishes its theoretical and empirical properties. Here, we generalize this variational distribution to apply to multiband coadded images.
@@ 3.2 Our variational family
Regier (2025) establishes conditions under which a variational distribution of this form mirrors the conditional independence assertions of the target distribution $p(z\mid x)$, in the sense of being a “minimally sufficient I-map.” This is a desirable property for a variational distribution because it implies that the variational distribution has the flexibility to model dependencies in the target distribution, but not enough flexibility to model dependencies that are known a priori to be absent in the target distribution.
@@ 3.3 Neural Posterior Estimation
During training, we additionally regularize by performing on-the-fly data augmentation. We follow standard practice in neural network performance optimization, which involves using validation loss to search for good network architectures and tuning parameters (Karpathy, 2019). Training is computationally intensive, requiring 22 hours on a single GPU. Inference is extremely fast, with data loading (moving images from disk to CPU memory) acting as the bottleneck rather than the neural network forward passes required to process that data. Appendix B provides further details of the proposed algorithm, including illustrations of the neural network architecture, projected runtimes for our methods on the complete LSST coadd image set, and discussion about storing probabilistic LSST catalogs.
@@ 4.1 Our primary benchmark: the LSST coadd catalog
The LSST pipeline constructs catalogs of objects in coadded images through a sequential multi-step process (Bosch et al., 2017).
@@ 4.2 Matching detections and true objects
To evaluate the accuracy of a catalog, we first match cataloged detections and true objects. The true objects available for matching are those exceeding 50 Jy in flux. For each of these objects, we identify the corresponding detection by selecting the nearest detected object within a one-pixel radius. A one-pixel radius is within the range of matching thresholds used in other publications: Sánchez et al. (2020) reports using a three-pixel radius for DC1 data and Portillo et al. (2017) reports using a half-pixel radius for SDSS data. Appendix C considers additional matching thresholds, namely three and five pixels; in short, the general trends in relative performance are not changed by the precise threshold used. Among detection-object pairs within the specified radius, we then find the optimal (distance-minimizing) bipartite matching using the Hungarian algorithm. The DC2 catalog provides a matching based on a different matching algorithm, but we exclusively use our matching procedure to ensure that the matching is done in the same way for both BLISS and the LSST pipeline.
@@ 4.6 Units of flux
The flux (or brightness) of an object directly influences the accuracy with which a method can detect that object. We quantify the brightness of an object in units of AB magnitude (Ivezić & the LSST Project Science Team, 2018). In this system, magnitude is defined as $2.5\log_{10}(F_{b}/F_{AB})$, where $F_{b}$ is the flux of an object in Jy (Jansky) and $F_{AB}$ is the zero point, which is set to 3631 Jy. Note that smaller magnitudes correspond to brighter sources.
@@ 4.8 Units of blending
We report some metrics stratified by blendedness, typically in the context of a minimum flux threshold. We adopt the high-level approach to quantifying blendedness of Bosch et al. (2017), in which blendedness is $1-\frac{I_{\textrm{child}}}{I_{\textrm{parent}}}$, where $I_{\textrm{child}}$ is the flux of the object in question (true or predicted, depending on context) and $I_{\textrm{parent}}$ is the overall flux in the vicinity of this object. The closer this value is to 1, the more blended the object is.
We define $I_{\textrm{child}}$ and $I_{\textrm{parent}}$ slightly differently than Bosch et al. (2017) because we alone have access to ground-truth catalogs by virtue of working with synthetic data. We take $I_{\textrm{child}}$ to be the true flux of the object in question (that is, without shot noise) and we take $I_{\textrm{parent}}$ to be the Gaussian-weighted sum of the true fluxes of neighboring objects. The covariance is isotropic with a standard deviation of 6 pixels (1.2 arc seconds), and we truncate beyond 30 pixels (6 arc seconds).
@@ 5 Experimental Results
In this section, we refer to our implementation of the proposed NPE method as the Bayesian Light Source Separator (BLISS). BLISS is a long-running software project (Liu et al., 2023; Hansen et al., 2022; Patel et al., 2025; Regier, 2025), which we have extended in this work to apply to the DC2 dataset. Our experiments use the benchmark methodology from Section 4 to assess the performance of BLISS in object detection, flux estimation, star/galaxy classification, and galaxy shape estimation.
@@ 5.1 Detection
To generalize the concepts of precision and recall to spatial data, in keeping with Portillo et al. (2017) and others, we define detection precision (purity) as the proportion of detections that are successfully matched with true objects and detection recall (completeness) as the proportion of objects that are successfully matched with detections. Figure 3 presents detection precision and recall as a function of magnitude for both BLISS and the LSST pipeline. The histogram in the top-right panel of Figure 3 shows the total number of true objects in each r-band magnitude bin. Compared to LSST, BLISS successfully identifies a larger fraction of sources in each of the magnitude bins. Within the r-band magnitude range of 23.9 to 25.6, the recall of LSST decreases from 93% to 58%, and the recall of BLISS exhibits a similar drop from 97% to 64%. The difference in recall between LSST and BLISS is largest in the second-to-last magnitude bin, where the difference in recall exceeds 10%. Even for the faintest sources (magnitude >> 25.6), which are the most difficult to detect, BLISS outperforms the LSST pipeline in terms of recall by 8%.
@@ 5.3 Star/galaxy classification
Classifying stars and galaxies in photometric surveys is essential for various scientific applications, including studies of galaxy formation and cosmology (Sevilla-Noarbe et al., 2018). In the field of weak lensing, identifying a pure sample of stars is crucial for fitting a model of the point spread function for use in inferring cosmic shear. Mistaking galaxies for stars could result in spurious correlations.
Star/galaxy separation is challenging because many light sources are faint and blended. Further, once convolved with the PSF, some galaxies appear scarcely larger than stars. The LSST pipeline does not use machine learning to separate stars and galaxies. Instead, it compares the fit of an extended “composite” model to that of a point source (Bosch et al., 2018)